OpenStudy (anonymous):

Brianna’s score on a national math assessment exceeded the scores of 95,000 of the 125,000 students who took the assessment. What was her percentile rank?

OpenStudy (anonymous):

Does anyone know how to do this problem?

OpenStudy (nurali):

\[\frac{ 95000 }{ 125000}*100\]

OpenStudy (anonymous):

where did you got 100 from?

OpenStudy (whpalmer4):

To convert to a percentage or percentile, multiply the fraction by 100. To convert a percentage to a fraction, divide by 100.

OpenStudy (whpalmer4):

27% = 27/100 = 0.27

OpenStudy (anonymous):

ok got it ty